# Existence of strong Lefschetz algebras with Chow polynomials as Hilbert series

### v1

@@ Abstract.
In this article, we study Chow polynomials of weakly ranked posets and prove the existence of Gorenstein algebras with the strong Lefschetz property such that their Hilbert–Poincaré series agrees with the Chow polynomial, providing evidence in support of a conjecture by Ferroni, Matherne and the second author. This allows us to show strong inequalities for the coefficients of Chow polynomials; we prove log-concavity for all posets of weak rank at most six and provide counterexamples to log-concavity for any higher rank. For ranked posets we recover an even stronger condition, showing that the differences between consecutive coefficients constitute a pure $O$-sequence.
@@
Artinian algebras, Chow polynomials, Poset theory, Monomial order ideals, O-sequences, Strong Lefschetz Algebras
@@ 1. Introduction
A central theme in the study of enumerative invariants of posets is the extent to which integer sequences satisfy strong regularity properties, such as unimodality, palindromicity, log-concavity, or the existence of combinatorial or algebraic models realizing them as Hilbert–Poincaré series. Chow polynomials of partially ordered sets, introduced recently by Ferroni, Matherne, and the second author, provide a unifying framework for encoding subtle enumerative and algebraic properties of posets [FMV24]. Their names come from the theory of combinatorial algebraic geometry for matroids; when the poset is a geometric lattice, they coincide with the Hilbert–Poincaré series of the Chow ring of the associated matroid. In this setting, Adiprasito, Huh and Katz proved that the Chow ring of a matroid satisfies a combinatorial analogue of the Kähler package [AHK18]. This collection of properties includes Poincaré duality and the strong Lefschetz property, which imply the palindromicity and unimodality of the corresponding Chow polynomials. For general posets there is not such a ring-theoretic interpretation, however, similar properties still hold. In particular, it is known that the coefficients of the characteristic Chow polynomial of any poset are non-negative, palindromic, and unimodal, and that they are also $\gamma$-positive for Cohen–Macaulay posets [FMV24]. An outstanding open problem is whether a ring with Poincaré duality and the strong Lefschetz property can be constructed for any poset so that its Hilbert–Poincaré series coincides with the Chow polynomial [FMV24, Section 4.6].
@@
The proof proceeds by giving an explicit combinatorial construction of a monomial order ideal whose graded pieces realize the successive differences of the Chow coefficients. A key ingredient is an analysis of Feichtner–-Yuzvinsky monomials, combined with symmetric chain decompositions of products of chains. Theorem 1.1 allows us to state the following corollary immediately.
@@
Let $P$ be a weakly ranked, finite, bounded poset of weak rank $n$. Then there exists a Gorenstein algebra with the Strong Lefschetz property whose Hilbert–Poincaré series coincides with the Chow polynomial of $P$.
Furthermore, we are able to show that this monomial ideal is pure, i.e., all its maximal monomials are of the same degree, if we assume that the poset is ranked. This is the second main result of our paper.
@@ 2.1. Properties of sequences
For two sequences corresponding to polynomials $p,q$ we refer to the monomial sequence of their product $pq$ as their convolution. It is well known that the following chains of (strict) implications hold:
@@ 2.2. Standard graded algebras
over a ring $R$ is said to be a standard graded algebra if $A_{0}\simeq R$ and $A$ is generated by the elements of $A_{1}$ as an algebra. In this paper, we will consider graded Artinian algebras over fields, that is, we may assume that we may express $A$ as $A=K[x_{1},x_{2},\ldots,x_{d}]/I$ and that
@@
An Artinian Gorenstein algebra $A$ satisfies the Weak Lefschetz property if it has a unimodal Hilbert–Poincaré series and if there exists an element $\ell\in A_{1}$ such that the map
@@
is an isomorphism for every $k$.
@@ 2.3. Chow polynomials
$H_{P}(t)=\sum_{s\geq 0}\sum_{\hat{0}=p_{0}<p_{1}<\ldots<p_{s}\leq\hat{1}}\prod_{i=1}^{s}\frac{t\left(t^{\rho(p_{i})-\rho(p_{i-1})-1}-1\right)}{t-1}.$
@@
For any poset $P$, the Chow polynomial is a monic polynomial of degree $\rho(P)-1$ with non-negative, palindromic, unimodal coefficients. Moreover, if $P$ is Cohen–Macaulay, then it is also $\gamma$-positive.
@@
In [FY04], Feichtner and Yuzvinsky defined a toric variety associated to any atomistic lattice and studied its Chow ring, giving an explicit presentation and a Gröbner basis for its ideal of relations. Later, Adiprasito, Huh and Katz showed that when $P$ is a geometric lattice the corresponding Chow ring is an Artinian Gorenstein algebra with the Strong Lefschetz property [AHK18]. These properties are easily seen to not hold for the Feichtner–Yuzvinsky Chow ring of generic atomistic lattices.
By comparing the Feichtner–Yuzvinsky Gröbner basis with Definition 2.3, it is clear that when $P$ is a geometric lattice, its Chow polynomial coincides with the Hilbert–Poincaré series of its Chow ring, see also [FMSV24, Theorem 1.5]. This, amongst other indicative numerical properties of the Chow polynomial led Ferroni, Matherne and the second author to conjecture in [FMV24, Section 4.6] the existence of a combinatorial ring generalizing the Chow ring for arbitrary weakly ranked posets exhibiting various strong properties, including Poincaré duality and the Strong Lefschetz property.
@@ 2.4. Monomial order ideals and (pure) $O$-sequences
Let $P$ be a poset of rank $n$ and let $H_{P}(t)=\sum_{i=0}^{n-1}h_{i}t^{i}$ its Chow polynomial. We call the sequence
@@
We justify this name by constructing a monomial order ideal $\operatorname{FY}$ with $h(P)$ as its $h$-vector. Consider the polynomial ring $\mathbb{Q}[x_{p}\mid p\in P\setminus\hat{0}$] and for a given chain $\hat{0}=p_{0}<p_{1}<\cdots<p_{s}\leq\hat{1}$ consider all monomials $m$ of the form
@@
where $1\leq\ell_{i}\leq d_{i}-1$ and $d_{i}=\rho(p_{i})-\rho(p_{i-1})$. We say that a monomial of this form is a Feichtner–Yuzvinsky monomial (or FY monomial for short). We denote by $\operatorname{FY}$ the family of all FY monomials and by $\operatorname{FY}^{k}$ the family of FY monomials of degree $k$. Clearly, $\operatorname{FY}$ is a, not necessarily pure, monomial order ideal. The following is a direct consequence of the definition of $H$ (the argument in [FY04, Corollary 1 and p. 526] works in general for posets).
@@
$\left\lfloor\frac{h_{1}+1}{2}\right\rfloor\leq h_{2}\leq\binom{h_{1}+1}{2}$
@@
Conversely, for any vector $(1,h_{1},h_{2})$ satisfying these properties we may take $h_{1}$ different variables $x_{1},x_{2},\ldots,x_{h_{1}}$ and choose $h_{2}$ degree two monomials to get a pure monomial order ideal. We start by choosing $\left\{x_{2i-1}x_{2_{i}}\right\}_{i=1}^{\left\lfloor h_{1}/2\right\rfloor}$ and then we add to the set the monomial $x_{h_{1}}^{2}$ if $h_{1}$ is odd. This subset has size exactly $\left\lfloor(h_{1}+1)/2\right\rfloor$ elements, and covers each of the $h_{1}$ degree one monomials. Thus, any choice for the remaining $h_{2}-\left\lfloor(h_{1}+1)/2\right\rfloor$ monomials gives a pure monomial order ideal. ∎
@@
  • $h_{0}\leq h_{1}\leq\ldots\leq h_{\lfloor e/2\rfloor}$.
@@ 2.5. $\operatorname{SI}$-sequences
$\Delta h=(1,h_{1}-1,h_{2}-h_{1},\ldots h_{\lfloor e/2\rfloor}-h_{\lfloor e/2\rfloor-1}).$
Sommerville proved in [Som27] that the $h$-vector of a simplicial polytope $P$ is unimodal and symmetric. McMullen [McM71] then conjectured that a vector is the $h$-vector of a simplicial polytope if and only if $\Delta h$, also known as the $g$-vector, is an $O$-sequence. This is commonly known as the $g$-conjecture for polytopes.
Later, McMullen defined in [McM93] the polytope algebra $\Pi_{P}$ for any simplicial polytope $P$. This is a standard graded algebra whose Hilbert series is equal to the $h$-vector of $P$ and with the Strong Lefschetz property. When taking a quotient of $\Pi_{P}$ by an ideal generated by any Lefschetz element, one obtains a standard graded algebra whose Hilbert–Poincaré series coincides with the $g$-vector, thus giving a proof for the necessity of the statement.
@@
  • $h$ is the Hilbert–Poincaré series of a Gorenstein algebra with the strong Lefschetz property.
Our goal now is to prove that the $\operatorname{FY}$ $h$-vector is a $\operatorname{SI}$-sequence. To do so, in analogy with the polytope case we define the $\operatorname{FY}$ $g$-vector of $P$ as $g=g(P)\coloneq\Delta h(P)$. By Theorem 2.5 the only thing we would need to show then is that $g$ is an $O$-sequence.
@@ 3. Symmetric chain decompositions
The goal of this section is to find a combinatorial interpretation of the $\operatorname{FY}$ $g$-vector of a poset. A symmetric chain decomposition of a poset $P$ of rank $n$ is a partition of $P$ such that each part is a saturated chain where if the first element is of rank $r$, the last element is of rank nrn-r. In [dBvETK51], the authors proved that if $P_{1}$ and $P_{2}$ are two posets admitting a symmetric chain decomposition, then so is the Cartesian product $P_{1}\times P_{2}$. Let $C_{n}$ denote the chain poset of rank $n$.
@@
Let $P=C_{1}\times C_{2}\times\cdots\times C_{s}$. Then $P$ admits a symmetric chain decomposition.
For a symmetric chain decomposition $S$ we will refer to the set of minimal elements of each saturated symmetric chain as the set of initial elements of $S$ denoted as $S_{\text{init}}$.
@@
Every element with rank 0 must be the start of a chain in the decomposition, thus, the number of initial elements with rank 0 is exactly $W_{0}$.
@@
We will now construct a specific symmetric chain decomposition for a product of chains, which exists by Corollary 3.1 and give an explicit description of its set of initial elements.
@@
$P=C_{r_{1}}\times C_{r_{2}}\times\ldots\times C_{r_{n}}$
@@
$S_{\text{init}}=\{(a_{1},\ldots,a_{n})\ |\ a_{s}\leq r_{s},\;a_{s}\leq\sum_{i=s+1}^{n}r_{i}-2a_{i}\}.$ (1)
@@
$C_{r_{1}}$$S^{\prime}_{i}$0112233445566011223344
The initial elements in this decomposition are exactly the ones where $0\leq a_{1}\leq\min(r_{1},\sum_{j=2}^{n}r_{j}-2a_{j})$ and $(a_{2},\ldots,a_{n})$ satisfies (1) for $P^{\prime}$. Taking the union of these decompositions over all $S^{\prime}_{i}$ we get exactly the set of initial elements described in the statement. ∎
@@
For any poset $P$ there exists a symmetric chain decomposition $S$ of $\operatorname{FY}$ such that the set of initial elements, which we denote by $\operatorname{SFY}$, coincides with the family of FY monomials of the form
@@
Let us build a symmetric chain decomposition of $\operatorname{FY}$. First, for a chain $\mathcal{C}=\hat{0}=p_{0}<p_{1}<p_{2}<\ldots<p_{s}<\hat{1}$ we define $\operatorname{FY}_{\mathcal{C}}$ to be the subposet of $\operatorname{FY}$ consisting of monomials with support either $\mathcal{C}$ or $\mathcal{C}\cup\{\hat{1}\}$. The bottom element of $\operatorname{FY}_{\mathcal{C}}$ is $\prod_{i}x_{p_{i}}$ with rank $s$, while its top element, $\prod_{i}x_{p_{i}}^{d_{i}}\ x_{\hat{1}}^{d_{\hat{1}}}$ has rank
@@
By Theorem 2.5, we only need to prove that $h$ forms a differential $O$-sequence. This follows by Theorem 3.5 and Theorem 4.1 as they imply that the $\operatorname{FY}$ $g$-vector is the $h$-vector of the monomial order ideal $\operatorname{SFY}$. The corollary follows by Theorem 2.15. ∎
@@
We show that any monomial that is maximal under divisibility has the correct degree, i.e., $\left\lfloor\frac{n-1}{2}\right\rfloor$. To do so, consider a maximal monomial $m$ and write $m=x_{p_{k}}^{\ell}\cdot m^{\prime}$, where $m^{\prime}$ is a monomial of degree $r$ with variables corresponding to elements that are strictly greater than the element $p_{k}$ of rank $k$. By contradiction, suppose that $\ell+r<\left\lfloor\frac{n-1}{2}\right\rfloor$. Clearly, $\ell=\min(n-k-2r,k-1)$, otherwise we could increase this exponent without violating the defining conditions, contradicting the maximality of $m$. We now distinguish two cases,
@@
Figure 2. Counter-example to pureness

### v2

# Existence of Kähler algebras with Chow polynomials as Hilbert series

@@ Abstract.
In this article, we study Chow polynomials of weakly ranked posets and prove the existence of Gorenstein algebras with the Kähler package such that their Hilbert–Poincaré series agrees with the Chow polynomial. Our statement provides evidence in support of a conjecture by Ferroni, Matherne and the second author about the existence of an algebra for every weakly ranked poset that generalizes the Feichtner–Yuzvinsky Chow ring for matroids. This allows us to prove strong inequalities for the coefficients of Chow polynomials; we prove log-concavity for all posets of weak rank at most six and provide counterexamples to log-concavity for any higher rank. For ranked posets we recover an even stronger condition, showing that the differences between consecutive coefficients constitute a pure $O$-sequence.
@@
Artinian algebras, Chow polynomials, Poset theory, Monomial order ideals, O-sequences, Kähler package
@@ 1. Introduction
A central theme in the study of enumerative invariants of posets is the extent to which integer sequences satisfy strong regularity properties, such as unimodality, palindromicity, log-concavity, or the existence of combinatorial or algebraic models realizing them as Hilbert–Poincaré series. A classical example of such a construction is due to Stanley [Sta06] and independently McMullen [McM93], who construct for every simplicial polytope $P$ a graded ring whose Hilbert–Poincaré series coincides with the $h$-vector of $P$. Chow polynomials of partially ordered sets, introduced recently by Ferroni, Matherne, and the second author, provide a unifying framework for encoding subtle enumerative and algebraic properties of posets [FMV24]. The name comes from the theory of combinatorial algebraic geometry for matroids; when the poset is a geometric lattice, it coincides with the Hilbert–Poincaré series of the Chow ring of the associated matroid, as defined by Feichtner and Yuzvinsky [FY04]. In this setting, Adiprasito, Huh and Katz proved that the Chow ring of a matroid satisfies a combinatorial analogue of the Kähler package [AHK18]. This collection of properties includes Poincaré duality, strong Lefschetz property and the Hodge–Riemann relations, which imply the palindromicity and unimodality of the corresponding Chow polynomials. For general posets there is not such a ring-theoretic interpretation, however similar properties still hold. In particular, it is known that the coefficients of the characteristic Chow polynomial of any poset are non-negative, palindromic, and unimodal, and that they are also $\gamma$-positive for Cohen–Macaulay posets [FMV24]. An outstanding open problem is whether a ring with Poincaré duality and the strong Lefschetz property can be constructed for any poset so that its Hilbert–Poincaré series coincides with the Chow polynomial [FMV24, Section 4.6].
@@
See Definitions 2.7 and 2.15 for the definition of $\operatorname{SI}$-sequence. This statement is equivalent to saying that $h(P)$ is the $h$-vector of a simplicial polytope, see also Theorem 2.16 below. The proof proceeds by giving an explicit combinatorial construction of a monomial order ideal whose graded pieces realize the successive differences of the Chow coefficients. A key ingredient is an analysis of Feichtner–-Yuzvinsky monomials, combined with symmetric chain decompositions of products of chains. Theorem 1.1 allows us to state the following corollary immediately.
@@
Let $P$ be a weakly ranked, finite, bounded poset of weak rank $n$. Then there exists a Gorenstein algebra with Kähler package whose Hilbert–Poincaré series coincides with the Chow polynomial of $P$.
Furthermore, we are able to show that if the poset is ranked, this monomial ideal is pure, i.e., all its maximal monomials are of the same degree. This is the second main result of our paper.
@@
Acknowldgments
The authors wish to thank Petter Brändén and Luis Ferroni for useful discussions.
@@ 2.1. Properties of sequences
For two sequences corresponding to polynomials $p,q$ we refer to the sequence of coefficients of their product $pq$ as their convolution. It is well known that the following chains of (strict) implications hold:
@@ 2.2. Standard graded algebras
over a ring $R$ is said to be a standard graded algebra if $A_{0}\simeq R$ and $A$ is generated by the elements of $A_{1}$ as an algebra. In this paper, we will consider graded Artinian algebras over fields, that is, we assume that we may express $A$ as $A=K[x_{1},x_{2},\ldots,x_{d}]/I$ and that
@@
An Artinian Gorenstein algebra $A=\bigoplus_{i=0}^{e}A_{i}$ satisfies the Weak Lefschetz property if there exists an element $\ell\in A_{1}$ such that the map
@@
is an isomorphism for every $k$. We refer to such an element $\ell$ as a Lefschetz element.
@@
Definition 2.3.
Let $A=\bigoplus_{i=0}^{e}A_{i}$ be an Artinian Gorenstein algebra over $\mathbb{R}$ with a Lefschetz element $\ell$. We say that $A$ satisfies the Hodge–Riemann relations if the bilinear maps defined as
$A_{k}\times A_{k}\to\mathbb{R}\quad(x,y)\mapsto(-1)^{k}d(xy\ell^{e-2k}),$
are positive definite when restricted to $\ker(\ell^{e-2k+1})\subset A_{k}$}.
An Artinian Gorenstein algebra over $\mathbb{R}$ has the Kähler package, if it satisfies Poincaré duality, the Strong Lefschetz property, and the Hodge-Riemann relations.
@@ 2.3. Chow polynomials
$H_{P}(t)=\sum_{s\geq 0}\sum_{\hat{0}=p_{0}<p_{1}<\cdots<p_{s}\leq\hat{1}}\prod_{i=1}^{s}\frac{t\left(t^{\rho(p_{i})-\rho(p_{i-1})-1}-1\right)}{t-1}.$
@@
For any poset $P$, the Chow polynomial is a monic polynomial of degree $\rho(P)-1$ with non-negative, palindromic and unimodal coefficients. Moreover, if $P$ is Cohen–Macaulay, then it is also $\gamma$-positive.
@@
In [FY04], Feichtner and Yuzvinsky defined a toric variety associated to any atomistic lattice and studied its Chow ring, giving an explicit presentation and a Gröbner basis for its ideal of relations. Later, Adiprasito, Huh and Katz showed that when $P$ is a geometric lattice the corresponding Chow ring is an Artinian Gorenstein algebra with the Kähler package [AHK18]. These properties are easily seen to not hold for the Feichtner–Yuzvinsky Chow ring of generic atomistic lattices.
By comparing the Feichtner–Yuzvinsky Gröbner basis with Definition 2.4, it is clear that when $P$ is a geometric lattice its Chow polynomial coincides with the Hilbert–Poincaré series of its Chow ring, see also [FMSV24, Theorem 1.5]. This, amongst other indicative numerical properties of the Chow polynomial led Ferroni, Matherne and the second author to conjecture in [FMV24, Section 4.6] the existence of a combinatorial ring generalizing the Chow ring for arbitrary weakly ranked posets exhibiting various strong properties, including Poincaré duality and the Strong Lefschetz property.
@@ 2.4. Monomial order ideals and (pure) $O$-sequences
Let $P$ be a poset of rank $n$ and let $H_{P}(t)=\sum_{i=0}^{n-1}h_{i}t^{i}$ be its Chow polynomial. We call the sequence
@@
We justify this name by constructing a monomial order ideal with $h(P)$ as its $h$-vector. Consider the polynomial ring $\mathbb{Q}[x_{p}\mid p\in P\setminus\hat{0}$] and all monomials $m$ of the form
@@
where $^{\hat{0}=p_{0}<p_{1}<\cdots<p_{s}\leq\hat{1}}$ is a chain of $P$, $1\leq\ell_{i}\leq d_{i}-1$ and $d_{i}=\rho(p_{i})-\rho(p_{i-1})$. We say that a monomial of this form is a Feichtner–Yuzvinsky monomial (or FY monomial for short). We denote by $\operatorname{FY}$ the family of all FY monomials and by $\operatorname{FY}^{k}$ the family of FY monomials of degree $k$. Clearly, $\operatorname{FY}$ is a, not necessarily pure, monomial order ideal. The following is a direct consequence of the definition of $H$ (the argument in [FY04, Corollary 1 and p. 526] works in general for posets).
@@
Conversely, for any vector $(1,h_{1},h_{2})$ satisfying these properties we may take $h_{1}$ different variables $x_{1},x_{2},\ldots,x_{h_{1}}$ and choose $h_{2}$ degree two monomials to get a pure monomial order ideal. We start by choosing $\left\{x_{2i-1}x_{2i}\right\}_{i=1}^{\left\lfloor h_{1}/2\right\rfloor}$ and then we add to the set the monomial $x_{h_{1}}^{2}$ if $h_{1}$ is odd. This subset has size exactly $\left\lfloor(h_{1}+1)/2\right\rfloor$ elements, and covers each of the $h_{1}$ degree one monomials. Thus, any choice for the remaining $h_{2}-\left\lfloor(h_{1}+1)/2\right\rfloor$ monomials gives a pure monomial order ideal. ∎
@@
  • $h_{0}\leq h_{1}\leq\cdots\leq h_{\lfloor e/2\rfloor}$.
@@ 2.5. $\operatorname{SI}$-sequences
$\Delta h=(1,h_{1}-1,h_{2}-h_{1},\ldots,h_{\lfloor e/2\rfloor}-h_{\lfloor e/2\rfloor-1}).$
Sommerville proved in [Som27] that the $h$-vector of a simplicial polytope $P$ is unimodal and palindromic. McMullen [McM71] then conjectured that a vector is the $h$-vector of a simplicial polytope if and only if $\Delta h$, also known as the $g$-vector, is an $O$-sequence. This is commonly known as the $g$-conjecture for polytopes.
Later, McMullen defined in [McM93] the polytope algebra $\Pi_{P}$ for any simplicial polytope $P$. This is a standard graded algebra with the Kähler package whose Hilbert series is equal to the $h$-vector of $P$. When taking a quotient of $\Pi_{P}$ by an ideal generated by any Lefschetz element, one obtains a standard graded algebra whose Hilbert–Poincaré series coincides with the $g$-vector, thus giving a proof for the necessity of the statement.
@@
  • $h$ is the Hilbert–Poincaré series of a Gorenstein algebra with the Strong Lefschetz property,
  • $h$ is the Hilbert–Poincaré series of a Gorenstein algebra with the Kähler package.
Remark 2.17.
In [Bre24, Problem 2.9], Brenti asks whether the coefficients of every monic, palindromic and real-rooted polynomial form a $\operatorname{SI}$-sequence. Solving the real-rootedness conjecture for Chow polynomials and Brenti’s conjecture would provide an independent proof of Theorem 1.1 for Cohen–Macaulay posets. We observe however that our result holds for a much larger class of posets.
Our goal now is to prove that the $\operatorname{FY}$ $h$-vector is a $\operatorname{SI}$-sequence. To do so, in analogy with the polytope case, we define the $\operatorname{FY}$ $g$-vector of $P$ as $g=g(P)\coloneq\Delta h(P)$. By Theorem 2.6 the only thing we would need to show then is that $g$ is an $O$-sequence.
@@ 3. Symmetric chain decompositions
The goal of this section is to find a combinatorial interpretation of the $\operatorname{FY}$ $g$-vector of a poset. A symmetric chain decomposition of a poset $P$ of rank $n$ is a partition of $P$ such that each part is a saturated chain, i.e., no element can be added between two consecutive elements without losing the property of being a chain, where if the first element is of rank $r$, the last element is of rank nrn-r. It is known that if $P_{1}$ and $P_{2}$ are two posets admitting a symmetric chain decomposition, then so is the Cartesian product $P_{1}\times P_{2}$ [dBvETK51]. Let $C_{n}$ denote the chain poset of rank $n$.
@@
The poset $P=C_{1}\times C_{2}\times\cdots\times C_{s}$ admits a symmetric chain decomposition.
For a symmetric chain decomposition $S$ we will refer to the set of minimal elements of each saturated symmetric chain as the set of initial elements of $S$, denoted as $S_{\text{init}}$.
@@
Every element of rank 0 must be the start of a chain in the decomposition, thus, the number of initial elements with rank 0 is exactly $W_{0}$.
@@
We will now construct a specific symmetric chain decomposition for a product of chains, which exists by Corollary 3.1, and give an explicit description of its set of initial elements.
$C_{r_{1}}$$S^{\prime}_{i}$0112233445566011223344
@@
$P=C_{r_{1}}\times C_{r_{2}}\times\cdots\times C_{r_{n}}$
@@
(1) $S_{\text{init}}=\big\{(a_{1},\ldots,a_{n})\ |\ a_{s}\leq r_{s},\;a_{s}\leq\sum_{i=s+1}^{n}r_{i}-2a_{i}\big\}.$
@@
The initial elements in this decomposition are exactly the ones where $(a_{2},\ldots,a_{n})$ satisfies (1) for $P^{\prime}$ and $0\leq a_{1}\leq\min(r_{1},\sum_{j=2}^{n}r_{j}-2a_{j})$}. Taking the union of these decompositions over all $S^{\prime}_{i}$ we get exactly the set of initial elements described in the statement. ∎
@@
For any weakly ranked poset $P$ there exists a symmetric chain decomposition $S$ of $\operatorname{FY}$ such that the set of initial elements, which we denote by $\operatorname{SFY}$, coincides with the family of FY monomials of the form
@@
Let us build a symmetric chain decomposition of $\operatorname{FY}$.
First, for a chain $\mathcal{C}=\{\hat{0}=p_{0}<p_{1}<p_{2}<\cdots<p_{s}<\hat{1}\}$ we define $\operatorname{FY}_{\mathcal{C}}$ to be the subposet of $\operatorname{FY}$ consisting of monomials with support either $\mathcal{C}$ or $\mathcal{C}\cup\{\hat{1}\}$. The bottom element of $\operatorname{FY}_{\mathcal{C}}$ is $\prod_{i}x_{p_{i}}$ with rank $s$, while its top element, $\prod_{i}x_{p_{i}}^{d_{i}}\ x_{\hat{1}}^{d_{\hat{1}}}$ has rank
@@
Remark 3.6.
We observe that when we apply Theorem 3.4, the resulting symmetric chain decomposition depends on an explicit choice of ordering the terms in the cartesian product. The choice we make in the proof above is needed for the set $\operatorname{SFY}$ to be a monomial order ideal, which is used in the proof of Theorem 1.1 below.
@@
The fact that $h$ forms a differential $O$-sequence follows from Theorem 3.5 and Theorem 4.1 as they imply that the $\operatorname{FY}$ $g$-vector is the $h$-vector of the monomial order ideal $\operatorname{SFY}$. The corollary follows by Theorem 2.16. ∎
@@
We show that any monomial that is maximal under divisibility has the correct degree, i.e., $\left\lfloor(n-1)/2\right\rfloor$. To do so, consider a maximal monomial $m$ and write $m=x_{p_{k}}^{\ell}\cdot m^{\prime}$, where $m^{\prime}$ is a monomial of degree $r$ with variables corresponding to elements that are strictly greater than the element $p_{k}$ of rank $k$. By contradiction, suppose that $\ell+r<\left\lfloor(n-1)/2\right\rfloor$. Clearly, $\ell=\min(n-k-2r,k-1)$, otherwise we could increase this exponent without violating the defining conditions, contradicting the maximality of $m$. We now distinguish two cases,
@@
Figure 2. Counterexample to pureness